a chemist needs a solution that is 75% acid but only has solutions that are 35% and 85% acid. If The chemist measures 120ml of the 85% solution, how many milliliters of the 35% solution should she add to make a 75% solution?
Answer by stanbon(75887)

a chemist needs a solution that is 75% acid but only has solutions that are 35% and 85% acid. If The chemist measures 120ml of the 85% solution, how many milliliters of the 35% solution should she add to make a 75% solution?
--------------------
Equation:
acid + acid = acid
0.35x + 0.85*120 = 0.75(x+120)
-----
Multiply thru by 100 to get:
35x + 85*120 = 75x + 75*120
40x = 10*120
x = 30ml (amt. of 35% solution needed for the mix)
==================
Cheers,
Stan H.
==================